HETAV PATEL

Grade 11,

Cot^2theeta=1+a^2 then cosectheeta+cot^3theetasectheeta=

Cot^2theeta=1+a^2 then cosectheeta+cot^3theetasectheeta=
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
I am writing theeta  = @ for the sake of convenience
 
cot^2 @ = 1 + a^2
 
cosec @ + cot^3 @ sec @ = 
 
 = cosec @ + cot^2 @ * cos @ / sin@ cos @ 
 
 = cosec @ [1 + cot^2 @]
 
 = sqrt (2 + a^2) [ 2 + a^2]
 
 = (2 + a^2)3/2
 
 
Regards
Arun (askIITians forum expert)